OpenStudy (anonymous):

sumation of n=1 going to infinity of ln (n/3n+1). Determine whether the series is convergent or divergent.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \ln (n/3n+1)\]

terenzreignz (terenzreignz):

This... \[\Large \sum_{n=1}^\infty \ln\left(\frac{n}{3n+1}\right)\] This?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Okay :) You know, the first thing you should check when you're testing convergence of series... is the underlying sequence, namely... \[\Large \ln\left(\frac{n}{3n+1}\right)\] What is ... \[\LARGE \lim_{n\rightarrow \infty}\quad\ln\left(\frac{n}{3n+1}\right)= \color{red}?\]

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

\[\ln (1/3)\]

terenzreignz (terenzreignz):

That's right :) And you know that for a series to be convergent, its underlying sequence must ALWAYS converge to ZERO right? This one doesn't... therefore...? ;)

OpenStudy (anonymous):

it is divergent.....but what test would you use to prove it? integral test? comparison test?

terenzreignz (terenzreignz):

Divergence test... basic :) The sequence does not converge to zero... that alone means the infinite series is divergent :)

terenzreignz (terenzreignz):

Your instructor might call it the "Limit Test" (Not the Limit COMPARISON test, ok?)

OpenStudy (anonymous):

oh that's right!!

terenzreignz (terenzreignz):

oh, you got it... great :D Nice job, by the way ;)

OpenStudy (anonymous):

haha thanks...ok so when would i know when to use this test?

terenzreignz (terenzreignz):

It's the easiest test to implement, but also possibly the least useful... I think your instructor is almost always going to give you a series whose underlying sequence does, in fact, converge to zero... btw, if the sequence converges to zero, it does NOT mean the series is convergent... It's easy to implement, anyway, so I suggest you do this limit test first... if possible, and who knows, you might get an instant point :)

OpenStudy (anonymous):

well if it converges to 0 it is inconclusive so i would have to use another test right?

terenzreignz (terenzreignz):

Yes. That's right :)

OpenStudy (anonymous):

ok got it thanks!

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

do you think you could help me out with the limit comparison test please?

terenzreignz (terenzreignz):

Sure :) You got a sample question?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (n ^{2} -5n)/ (n ^{3}+n+1)\]

terenzreignz (terenzreignz):

Interesting :) Okay, to use the Limit-Comparison test, much like with Direct Comparison, you need an intuition... what do you suspect this series to be... convergent, or divergent?

OpenStudy (anonymous):

well i would compare it to 1/n and that diverges, so divergent?

terenzreignz (terenzreignz):

That's right :) So, you know the gist of the Limit-Comparison test?

terenzreignz (terenzreignz):

You put both sequences, one on top of the other (LOL) and get the limit as n goes to infinity... \[\LARGE \lim_{n\rightarrow \infty}\frac{a_n}{b_n}\] If it goes to some number greater than 0, then either both converge or both diverge.

terenzreignz (terenzreignz):

If it goes to 0, then if a_n converges, then so does b_n If it goes to infinity, then if b_n converges, then so does a_n

OpenStudy (anonymous):

if it goes to some number greater than zero... wouldn't the series diverge? My Professor said it congerves...but the p-series tests says it diverges when you compare it to 1/n...that is what is confusing me

terenzreignz (terenzreignz):

No, because we're not taking the limit of just one sequence, we're taking the limit of the ratio of two sequences... so that if the limit of the ratio is some positive number, then their convergences are tied to each other (ie, either both converge or both diverge) the proof is a bit extensive...

OpenStudy (anonymous):

ok so if it's greater than zero, it converges? and if it's less than zero it would diverge?

terenzreignz (terenzreignz):

A requirement for the limit comparison test is that both series are positive... so it's kinda impossible for the limit to be negative :3

OpenStudy (anonymous):

ohhhh ok....so when would a series diverge?

terenzreignz (terenzreignz):

And no, if the limit is greater than zero, then either BOTH series converge or BOTH diverge...

terenzreignz (terenzreignz):

Here's how it works... you have a series... you suspect it's convergent (divergent). So you compare it to a known convergent (divergent) series, and use the limit comparison test... If the limit is a positive number, then your original series mus have been convergent (divergent), since a positive limit implies both series to be either convergent or divergent.

OpenStudy (anonymous):

ok i think i got it. Will you be here long? I think i'm going to need help in other tests. I'm reviewing for my test this wednesday

terenzreignz (terenzreignz):

Probably not... sorry... I'm sure there are others here that can help you out :)

OpenStudy (anonymous):

i hope so. Thank you for your help!